an isosceles triangle has an area of 200 square units, and the altitude to the base measures 10 units. what is the perimeter of the triangle?
Answer by checkley71(8403)

1/2*b*h=200 or b/2*10=200 or 5b=200 or b=40 or now we have a two right triangles with the two sides 10 & 20 then the hyp is 10~2+20~2=hyp~2 or
100+400=hyp~2 or 500=hyp~2 or hyp=sqrt500 or hyp=22.36 then we have a perimeter of two sides of 22.36+22.36 and the base of 40 for a total perimeter of 2*22.36+40=84.72